OpenStudy (anonymous):

the sum of the terms of the arithmetic sequence. 1, 7, 13, 19, . . ., 109 n = 19

OpenStudy (solomonzelman):

you add 6 every time, and thus 109 = 108+1=1 + (6 × 18) so 109 is the 18th term. So, you can say now, S(n) = ½ (1+109) × 18

OpenStudy (solomonzelman):

Need more help ?

OpenStudy (solomonzelman):

my bad 109 is the 19th term

OpenStudy (solomonzelman):

because you skip 18 times

OpenStudy (solomonzelman):

S(n) = ½ (1+109) × 19 (not 18, but 19) Like this.

OpenStudy (anonymous):

\[t_{n}=a+(n-1)d,109=1+(n-1)6,109-1=6(n-1),n-1=\frac{ 108 }{ 6 }=18\] n=18+1=19 \[s _{n}=\frac{ n \left( a+t _{n} \right) }{ 2 },s _{19}=\frac{ 19\left( 1+109 \right) }{ 2 }=\frac{ 19\times110 }{ 2 }=?\]